OpenStudy (anonymous):

gais xD, can you give the answer to this and explain? Two numbers are in the ratio 4:7. If the difference is 39, what are the two numbers?

OpenStudy (anonymous):

let the numbers be 4x and 7x.. den 7x-4x = 39 3x = 39 x = 13 numbers wud be.. 4*13 = 52 and 7*13 = 91

OpenStudy (anonymous):

k tnx bro

OpenStudy (anonymous):

wait how did you get 13 (sorry i don't know much on math)

OpenStudy (anonymous):

since 3x = 39, divide both sides by 3 to get x isolated. therefore 3x/3 = 39/3 3x/3 = x and 39/3 = 13 therefore x=13

OpenStudy (anonymous):

k bro i get it nao NVM tnx